OpenStudy (anonymous):

Find the distance between the points (-3, 11) and (5, 5). 10 2 2

OpenStudy (experimentx):

10

OpenStudy (callisto):

use distance formula distance \[=\sqrt{[(-3)-5]^2 + (11-5)^2}\] \[=\sqrt{(64)+(36)}\] \[=\sqrt{100}\] =10